A transformer connected to a 120-V (rms) ac line is to supply 13,000 V (rms) for a neon sign. To reduce shock hazard, a fuse is to be inserted in the primary circuit; the fuse is to blow when the rms current in the secondary circuit exceeds 8.50 mA. What current rating should the fuse in the primary circuit have? (Points : 20)
Answer #1

we know,

POWER DELIVERD TO PRIMARTY CIRCUIT = POWER DELIVERED TO SECONDARY CKT

V1*I1 = V2*I2

I1 = I2*(V2/V1)

= 8.5*10^-3*(13000/120)

= 0.92 A